OpenStudy (anonymous):

Which of the following is not an example of a parent function? A. y = x B. y=2x^3 C. y=x^x D. y=2^x

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@cj49

OpenStudy (anonymous):

its fine

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@precal

OpenStudy (precal):

parent functions don't have any transformations example y=x y=x^2 y=x^3 it is your basic function

OpenStudy (anonymous):

here we go again

OpenStudy (anonymous):

???

OpenStudy (anonymous):

well i know its not b

OpenStudy (anonymous):

why not?

OpenStudy (anonymous):

because of the 2 before the x

OpenStudy (precal):

true

OpenStudy (anonymous):

i dont think its d either

OpenStudy (anonymous):

it stays which one is NOT right?

OpenStudy (anonymous):

woops

OpenStudy (anonymous):

hahaha

OpenStudy (anonymous):

sorry